# A geometric perspective on Algebraic Quantum Field Theory

### v1

@@ 2.4.1 Scalar free fields in Algebraic Quantum Field Theory
This representation extends by the same formula to a (anti-)unitary representation of the proper Poincaré group $\mathcal{P}_{+}$, [Va85, Thm. 9.10]. We can define the real subspace
${\sf H}(\mathcal{O})=\overline{\{\widehat{f}|_{\Omega_{m}}:f\in\mathcal{C}_{0}^{\infty}(M),\mathop{{\rm supp}}\nolimits f\,\subset\mathcal{O}\}}$
with $\mathcal{O}\subset M$ open bounded region. This is the set of one-particle states localized in $\mathcal{O}$. Note that the following properties hold
@@
  • Reeh–Schlieder: $\overline{{\sf H}(\mathcal{O})+i{\sf H}(\mathcal{O})}=\mathcal{H}$ for all $\mathcal{O}\subset M$.
@@
The free field is the second quantization net of von Neumann algebras on the Fock space. It is constructed as follows. Given the one-particle Hilbert space $\mathcal{H}$, we can define the Fock space
@@
where $\mathcal{H}^{\otimes_{s}^{n}}$ is the symmetrized $n$-fold tensor product of the one-particle Hilbert space $\mathcal{H}$.
For every vector $\xi\in\mathcal{H}$ the associated Weyl unitary $\mathrm{w}(\xi)$ on the Fock space is defined by the commutation relations
@@
Note that the map $\mathcal{H}\ni h\mapsto\mathrm{w}(h)\in\mathcal{F}(\mathcal{H})$ is strongly continuous. Given a real subspace $H\subset\mathcal{H}$, we can associate a the von Neumann algebra acting on the Fock space as follows
@@
  • Poincaré covariance and positivity of the energy: there exists a (anti-)unitary representation $\widetilde{U}$ of $\mathcal{P}_{+}$ on $\mathcal{F}(\mathcal{H})$ such that $\widetilde{U}(g)\Omega=\Omega$ and for $g\in\mathcal{P}_{+}$, we have $\widetilde{U}(g)\mathcal{A}(\mathcal{O})\widetilde{U}(g)^{*}=\mathcal{A}(gO)$. Furthermore the joint spectrum of the tranlsation generators is contained in the forward lightcone $C$.
@@
We recall that, given a von Neumann algebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$ with a cyclic and separating vector $\Omega\in\mathcal{H}$, one can define the Tomita operator $S_{\mathcal{A},\Omega}$ as the closure of the anti-linear involution
The polar decomposition $S_{\mathcal{A},\Omega}=J_{\mathcal{A},\Omega}\Delta_{\mathcal{A},\Omega}^{1/2}$ defines the self-adjoint modular operator $\Delta$ and the anti-unitary modular conjugation $J_{\mathcal{A},\Omega}$ that satisfy $J_{\mathcal{A},\Omega}\Delta_{\mathcal{A},\Omega}J_{\mathcal{A},\Omega}=\Delta_{\mathcal{A},\Omega}^{-1}$.
@@
$\widetilde{U}(\tau_{h_{R}})=J_{\mathcal{A}(W_{R}),\Omega}.$
We remark that this is weaker than the PCT Theorem. The latter ensures an antiunitary implementation of the space, time and charge reflection on the Hilbert space (not only of the wedge coordinates reflection). In even spacetime dimension, the space and time reflection differs from $\tau_{W_{R}}$ by $\pi$-rotations, hence the two statements are equivalent, see [BY01].
@@ 2.4.2 Nets of standard subspaces
The polar decomposition $J_{\sf H}\Delta_{\sf H}^{\frac{1}{2}}$ of this operator defines an antiunitary involution $J_{\sf H}$ (a conjugation) and the modular self-adjoiont operator $\Delta_{\sf H}$. We then have
@@
and the modular relations
@@
We remark that given a von Neumann algebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$ with a cyclic and separating vector $\Omega\in\mathcal{H}$ then one can define the standard subspace ${\sf H}_{\mathcal{A}}=\overline{\mathcal{A}_{\text{s}a}\Omega}$, where $\mathcal{A}_{\text{s}a}=\{a\in\mathcal{A}:a=a^{*}\}$. It is easy to see that $S_{{\sf H}_{\mathcal{A}}}=S_{\mathcal{A},\Omega}$, namely the standard subspace ${\sf H}_{\mathcal{A}}$ contains all the information about the Tomita operator of the von Neumann algebra $\mathcal{A}$ with respect to the vector state $\Omega$.
One can completely reconstruct the standard subspace from the Tomita operators as
${\sf H}=\mathop{{\rm Fix}}\nolimits(J_{\sf H}\Delta_{\sf H}^{\frac{1}{2}})=\ker(1-J_{\sf H}\Delta_{\sf H}^{\frac{1}{2}})$ (14)
([Lo08, Thm. 3.4]). This construction leads to a one-to-one correspondence between couples $(\Delta,J)$ satisfying the modular relation:
@@
and standard subspaces ${\sf H}=\mathop{{\rm Fix}}\nolimits(J\Delta^{\frac{1}{2}})$, see [Lo08, Prop. 3.2].
@@
  • Covariance property:([Mo18, Lem. 2.2]). Let ${\sf H}\subset\mathcal{H}$ be a standard subspace and $U\in\mathop{{\rm AU}}\nolimits(\mathcal{H})$ be a unitary or anti-unitary operator. Then $U{\sf H}$ is also standard and $U\Delta_{\sf H}U^{*}=\Delta_{U{\sf H}}^{\epsilon(U)}$ and $UJ_{\sf H}U^{*}=J_{U{\sf H}}$, where $\epsilon(U)=1$ if $U$ is unitary and $\epsilon(U)=-1$ if it is anti-unitary.
@@
  • Inclusion property ([Lo08, Thms. 3.15, 3.17], [BGL02, Thm. 3.2]): The results we indicate here relate an inclusion property of standard subspaces and positivity of the generator of a related one-parameter group.
    Let ${\sf H}\subset\mathcal{H}$ be a standard subspace and $U(t)=e^{itP}$ be a unitary one-parameter group on $\mathcal{H}$ with generator $P$.
@@
Let $G$ be a connected Lie group with trivial center, such that $\mathcal{E}({\mathfrak{g}})\neq\{0\}$. We fix an Euler element and $h\in{\mathfrak{g}}$ and assume that $\tau_{h}$ integrates to an involution on $G$ and let $W_{0}=(h,\tau_{h})$. Let $C$ be an $\mathop{{\rm Ad}}\nolimits^{\epsilon}(G_{\tau_{h}})$-invariant pointed convex cone in the Lie algebra ${\mathfrak{g}}$ of $G$. Let $(U,\mathcal{H})$ be a unitary representation of $G$ and
@@
be a map, also called a net of standard subspaces. We consider the following properties:
@@
    Locality: If $W\in\mathcal{W}_{+}$ is such that $W^{\prime}\in\mathcal{W}_{+}$, then ${\sf N}(W^{\prime})\subset{\sf N}(W)^{\prime}.$
@@
    Bisognano–Wichmann (BW) property: $U(\exp(th))=\Delta_{{\sf N}(W)}^{-it/2\pi}$ for $t\in{\mathbb{R}}$, $W\in\mathcal{W}_{+}$.
@@
    Modular reflection: $U(\tau_{W})=J_{{\sf N}(W)}$, for $W=(h_{W},\tau_{W})\in\mathcal{W}_{+}$
It is a consequence of (HK8) that Haag duality holds for $W,W^{\prime}\in\mathcal{W}$.
@@
Note that it does not require the Bisognano–Wichmann property, but it holds whenever the latter is satisfied.
We remark that by identifying the abstract wedges with concrete wedge regions of Minkowski space, the net of standard subspaces on wedges of the free scalar fields given by (9) satisfies all the previous assumptions. The deep relation between wedge subspaces, wedge regions and the symmetry group representation has been made more explicit by the Brunetti–Guido–Longo construction as will be explained in the next section.
@@ 2.4.3 The Brunetti–Guido–Longo (BGL) net
satisfy the commutation relation (15), hence $\Delta_{W}=\exp(2\pi i\partial U(h_{W}))$ and $J_{W}=U(\tau_{h_{W}})$ canonically define a the wedge standard subspace
@@
as we have seen in the previous section. Based on these remarks R. Brunetti, D. Guido and R. Longo in [BGL02] provided a construction of the one-particle net of standard subspaces of the free Poincaré and conformal covariant theories on Minkowski space, de Sitter space and the chiral circle on wedge regions. This construction has been generalized to our abstract set of wedges for a ${\mathbb{Z}}_{2}$-graded Lie group supporting Euler elements in [MN21]. The anti-de Sitter space can be included in the picture and it is treated in [NÓØ21, NÓ23], see also [Re00].
Let $G$ be a connected Lie group with trivial center, such that $\mathcal{E}({\mathfrak{g}})\neq\{0\}$. We fix an Euler element and $h\in{\mathfrak{g}}$ and $\tau_{h}$ integrates to an involution on $G$. Let $(U,G)$ be an (anti-)unitary representation of $G_{\tau_{h}}$, $W=(h,\tau_{h})\in\mathcal{G}_{E}$ an Euler wedge, and consider the couple of operators
@@
[MN21, Thm 4.12, Prop. 4.16] Let $(U,\mathcal{H})$ be an (anti-)unitary C-positive representation of $G_{\tau}$ where $\tau$ is an Euler involution. Then the BGL net
${\sf H}_{U}^{\rm BGL}\colon\mathcal{G}(G_{\tau})\to\mathop{{\rm Stand}}\nolimits(\mathcal{H})$
@@
The statement of this theorem can be more general as it is presented in [MN21]. Given a ${\mathbb{Z}}_{2}$-graded Lie group $G_{\sigma}=G\rtimes\{e,\sigma\}$ with $\sigma$ an involution on $G$, then one can define more general abstract wedges as
@@
without referring to Euler elements, satisfying analogous properties to Euler wedges [MN21, Sect. 2] and the BGL net is still well defined. The reason why we will mainly deal with Euler element will become clear in Sect. 3.2.
We have introduced an axiomatic framework for nets of standard subspaces on wedges. This picture includes the well known models from AQFT and produces new examples. Indeed, all three graded Lie algebras support unitary representations of the $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$-group that, given $h\in\mathcal{E}({\mathfrak{g}})$, extend to $G_{\tau_{h}}$. Up to coupling conjugate representations, $U$ extends to an (anti-)unitary representation of $G_{\tau_{h}}$ and one can construct the BGL net of standard subspaces on abstract wedge regions. Hermitian simple Lie algebras have a proper positive cone and their unitary positive energy representations are studied in [MdR07]. Once a net of standard subspaces on abstract wedge regions is constructed the second quantization net of von Neumann algebras is canonically provided by relations (10) and (11). Furthermore, Theorem 22 also works if $C=C_{U}=\{0\}$. Then positivity of the energy is trivially satisfied as well as the isotony property. We stress that if the group is not centerfree the locality property has to be rewritten using twisted central complements. This has been investigated in [MN21, Sect. 4.2.2]. In this case a proper second quantization procedure respecting a twisted locality conditions has still to be established. Recent results on second quantizations are contained in [CSL23].
@@
Given a unitary representation $(U,\mathcal{H})$ of a connected a Lie group $G$ and a homogeneous space $M=G/H$, we are interested in families $({\sf H}(\mathcal{O}))_{\mathcal{O}\subseteq M}$ of closed real subspaces of $\mathcal{H}$, indexed by open subsets $\mathcal{O}\subseteq M$, the so-called nets of real subspaces on $M$.
@@
    Bisognano–Wichmann property: There exists an open subset $W\subseteq M$ (called a wedge region), such that ${\sf H}(W)$ is standard with modular operator $\Delta_{{\sf H}(W)}=e^{2\pi i\partial U(h)}$ for some $h\in{\mathfrak{g}}$.
Nets satisfying (Iso), (Cov), (RS), (BW) on non-compactly causal symmetric spaces have been constructed in [FNÓ23], and on compactly causal spaces in [NÓ23] using the language of distribution vectors. Given the $G$-orbit of the wedge defined, one can define the maximal net on bounded regions by wedge intersections ${\sf H}^{\text{max}}(\mathcal{O})=\bigcap\{{\sf H}(gW):\mathcal{O}\subset gW,g\in G\}$, this has been studied in [MN24, Sect. 2.2.4]. We remark that locality is not present among the list, since it is to be investigated [MN25].

### v2

@@ 2.4.1 Scalar free fields in Algebraic Quantum Field Theory
The free scalar mass $m\geq 0$ Quantum Field Theory on Minkowksi spacetime can be constructed as follows. Let $U$ be the unitary, scalar, mass $m$ positive energy representation of the proper Poincaré group $\mathcal{P}_{+}^{\uparrow}={\mathbb{R}}^{1,3}\rtimes\mathop{{\rm SO}}\nolimits(1,3)_{e}$. The Hilbert space is $\mathcal{H}=L^{2}(\Omega_{m},d\Omega_{m})$, where
$\Omega_{m}=~{}\{p\in~{}M:~{}p^{2}=m^{2},p_{0}>0\}$
@@
$\widehat{\mathcal{S}}(M)=\{\widehat{f}|_{\Omega_{m}}:f\in\mathcal{S}(M)\}$
defines a complex linear dense subset of $\mathcal{H}$. On this Hilbert space the unitary scalar representation acts by
@@
Positivity of the energy means that the joint spectrum of the translation generators is contained in the forward lightcone $C$. This representation extends by the same formula to a (anti-)unitary representation of the proper Poincaré group $\mathcal{P}_{+}$, [Va85, Thm. 9.10]. To any open bounded $\mathcal{O}\subset M$ we can associate the real subspace
${\sf H}(\mathcal{O})=\overline{\{\widehat{f}|_{\Omega_{m}}:f\in\mathcal{C}_{0}^{\infty}(M),\mathop{{\rm supp}}\nolimits f\,\subset\mathcal{O}\}}.$
This is the set of one-particle states localized in $\mathcal{O}$. Note that the following properties hold
@@
  • Reeh–Schlieder: $\overline{{\sf H}(\mathcal{O})+i{\sf H}(\mathcal{O})}=\mathcal{H}$ for all $\mathcal{O}\neq\emptyset$.
@@
The free field is the second quantization net of von Neumann algebras on the Fock Hilbert space. It is constructed as follows. Given the one-particle Hilbert space $\mathcal{H}$, we can define the Fock space
@@
where $\mathcal{H}^{\otimes_{s}^{n}}$ is the symmetrized $n$-fold tensor product of the one-particle Hilbert space $\mathcal{H}$. For every vector $\xi\in\mathcal{H}$ the associated Weyl unitary $\mathrm{w}(\xi)$ on the Fock space is defined by the commutation relations
@@
Then the map $\mathcal{H}\ni h\mapsto\mathrm{w}(h)\in\mathcal{F}(\mathcal{H})$ is strongly continuous. Further details in [BR87]. Given a real subspace $H\subset\mathcal{H}$, we can associate a von Neumann algebra acting on the Fock space as follows
@@
  • Poincaré covariance and positivity of the energy: there exists a (anti-)unitary representation $\widetilde{U}$ of $\mathcal{P}_{+}$ on $\mathcal{F}(\mathcal{H})$ such that $\widetilde{U}(g)\Omega=\Omega$ and for $g\in\mathcal{P}_{+}$, we have $\widetilde{U}(g)\mathcal{A}(\mathcal{O})\widetilde{U}(g)^{*}=\mathcal{A}(gO)$. Furthermore the joint spectrum of the translation generators is contained in the forward lightcone $C$.
@@
$\widetilde{U}(\tau_{h_{R}})=J_{\mathcal{A}(W_{R}),\Omega}.$ (15)
In even spacetime dimension the PCT Theorem ensures an antiunitary implementation of the space, time and charge reflection on the Hilbert space. Since the space and time reflection differs from $\tau_{h_{R}}$ by $\pi$-rotations, the modular reflection property is equivalent to the PCT-theorem, see [Mu01, BY01]. In odd spacetime dimension, the (15) is considered as the PCT theorem (cf. [Mu10, BY01]).
@@ 2.4.2 Nets of standard subspaces
The polar decomposition $S_{\sf H}=J_{\sf H}\Delta_{\sf H}^{\frac{1}{2}}$ defines an antiunitary involution $J_{\sf H}$, called the modular conjugation of ${\sf H}$, and a positive self-adjoiont operator $\Delta_{\sf H}$, called the modular operator of ${\sf H}$. Then we have that
@@
and the following modular relation holds, namely
@@
We remark that given a von Neumann algebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$ with a cyclic and separating vector $\Omega\in\mathcal{H}$ then one can define the standard subspace ${\sf H}_{\mathcal{A}}=\overline{\mathcal{A}_{\text{s}a}\Omega}$, where $\mathcal{A}_{\text{s}a}=\{a\in\mathcal{A}:a=a^{*}\}$. By direct inspection one can see that $S_{{\sf H}_{\mathcal{A}}}=S_{\mathcal{A},\Omega}$, namely the standard subspace ${\sf H}_{\mathcal{A}}$ contains all the information about the Tomita operator of the von Neumann algebra $\mathcal{A}$ with respect to the vector state $\Omega$.
One can completely reconstruct the standard subspaces from the related Tomita operators as
${\sf H}=\mathop{{\rm Fix}}\nolimits(S_{\sf H})=\ker(1-S_{\sf H})$ (17)
([Lo08, Thm. 3.4]). This construction leads to a one-to-one correspondence between couples $(\Delta,J)$ where $J$ is an antiunitary operator and $\Delta$ a selfadjoint positive operator, satisfying the modular relation:
@@
and standard subspaces ${\sf H}=\mathop{{\rm Fix}}\nolimits(J\Delta^{\frac{1}{2}})$, see [Lo08, Prop. 3.2]. Note that (18) is equivalent to $J\Delta J=\Delta^{-1}$.
@@
  • Covariance property:([Mo18, Lem. 2.2]). Let ${\sf H}\subset\mathcal{H}$ be a standard subspace and $U\in\mathop{{\rm AU}}\nolimits(\mathcal{H})$ be a unitary or antiunitary operator. Then $U{\sf H}$ is also standard and $U\Delta_{\sf H}U^{*}=\Delta_{U{\sf H}}^{\epsilon(U)}$ and $UJ_{\sf H}U^{*}=J_{U{\sf H}}$, where $\epsilon(U)=1$ if $U$ is unitary and $\epsilon(U)=-1$ if it is antiunitary.
@@
  • Inclusion properties ([Lo08, Thms. 3.15, 3.17], [BGL02, Thm. 3.2]): Let ${\sf H}\subset\mathcal{H}$ be a standard subspace and $U(t)=e^{itP}$ be a unitary one-parameter group on $\mathcal{H}$ with generator $P$.
@@
Let $G$ be a connected Lie group with trivial center, such that $\mathcal{E}({\mathfrak{g}})\neq\emptyset$. We fix an Euler element $h\in{\mathfrak{g}}$, then $\tau_{h}$ integrates to an involution on $G$. Let $W_{0}=(h,\tau_{h})$ be an Euler wedge and $C$ be an $\mathop{{\rm Ad}}\nolimits^{\epsilon}(G_{\tau_{h}})$-invariant pointed convex cone in the Lie algebra ${\mathfrak{g}}$ of $G$. Let $(U,\mathcal{H})$ be a unitary representation of $G$ and
@@
be a map called a net of standard subspaces on abstract wedges. We consider the following properties:
@@
    Locality: If $W_{1},W_{2}\in\mathcal{W}_{+}$ is such that $W_{1}\leq W_{2}^{\prime}$, then ${\sf N}(W_{1})\subseteq{\sf N}(W_{2})^{\prime}.$
@@
    Bisognano–Wichmann (BW) property: for every $W=(h_{W},\tau_{W})\in\mathcal{W}_{+}$
    $U(\exp(th_{W}))=\Delta_{{\sf N}(W)}^{-it/2\pi},\qquad\forall t\in{\mathbb{R}},$
@@
    Modular reflection: $U(\tau_{W})=J_{{\sf N}(W)}$ for every $W=(h_{W},\tau_{W})\in\mathcal{W}$.
We remark that if $C=\{0\}$, there are no non-trivial wedge inclusions, then the isotony property trivializes and the locality property reduces to the condition ${\sf H}(W^{\prime})\subseteq{\sf H}(W)^{\prime}$ when $W,W^{\prime}\in\mathcal{W}_{+}$.
@@
Note that the wedge modular covariance property does not require the Bisognano–Wichmann property, but it holds whenever the latter is satisfied.
By identifying the abstract wedges with concrete wedge regions of Minkowski space, the net of standard subspaces on wedges of the free scalar field given by (11) satisfies all the previous assumptions.
Lastly, we remark that an analogous set of axioms for nets of von Neumann algebras on abstract wedges that fits with second quantization can be provided ([MN21, MNO25]).
@@ 2.4.3 The Brunetti–Guido–Longo (BGL) net
satisfy the commutation relation (18), hence $\Delta_{W}=\exp(2\pi i\partial U(h_{W}))$ and $J_{W}=U(\tau_{h_{W}})$ define a the wedge standard subspace
@@
as we have seen in the previous section. Based on these remarks R. Brunetti, D. Guido and R. Longo in [BGL02] provided a construction of the one-particle net of standard subspaces of the free Poincaré and conformal covariant theories on Minkowski space, Lorentz covariant theory on de Sitter space and Möbius covariant theory on the chiral circle, on the corresponding index sets of wedge regions (see also [GLW98, Lo08]). This construction has been generalized to our abstract set of wedges for a ${\mathbb{Z}}_{2}$-graded Lie group supporting Euler elements in [MN21]. The anti-de Sitter space can be included in the picture and it is treated in [NÓØ21, NÓ23], see also [Re00].
Let $G$ be a connected Lie group with trivial center, such that $\mathcal{E}({\mathfrak{g}})\neq\emptyset$. We fix an Euler element $h\in{\mathfrak{g}}$ and $\tau_{h}$ integrates to an involution on $G$. Let $(U,\mathcal{H})$ be an (anti-)unitary representation of $G_{\tau_{h}}$, $W=(h,\tau_{h})\in\mathcal{G}_{E}$ be an Euler wedge, and consider the couple of operators
@@
[MN21, Thm 4.12, Prop. 4.16] Let $(U,\mathcal{H})$ be an (anti-)unitary C-positive representation of $G_{\tau_{h}}$ where $h\in\mathcal{E}({\mathfrak{g}})$. Then the BGL net
${\sf H}_{U}^{\rm BGL}\colon\mathcal{G}(G_{\tau_{h}})\to\mathop{{\rm Stand}}\nolimits(\mathcal{H})$
@@
Theorem 22 can be stated in a more general framework of abstract non-Euler wedges as in [MN21]. Given a ${\mathbb{Z}}_{2}$-graded Lie group $G_{\sigma}=G\rtimes\{e,\sigma\}$ with $\sigma$ an involution on $G$, then one can define more general notion for the set of abstract wedges without referring to Euler elements as follows:
@@
This set satisfies analogous properties to the set of Euler wedges and the BGL net is still well defined on $\mathcal{G}$ given an (anti-)unitary representation of $G_{\sigma}$ [MN21, Sect. 2]. In Section 3.2 we will explain why Euler elements, hence Euler wedges, are the natural choices for an index set of abstract wedges.
We have introduced an axiomatic framework for nets of standard subspaces on wedges. This picture includes the well known models from AQFT and produces new examples. Given any three graded Lie algebra ${\mathfrak{g}}$ and $h\in\mathcal{E}({\mathfrak{g}})$, the group $G=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ supports unitary representations that extend (anti-)unitarily to $G_{\tau_{h}}$ (up to properly coupling representations). Consider $\mathcal{G}_{E}(G_{\tau_{h}})$ the set of Euler wedges and $\mathop{{\rm Ad}}\nolimits^{\varepsilon}(G_{\tau_{h}})$-invariant pointed cone $C$ in ${\mathfrak{g}}$ defining the inclusion rules of Euler wedges. Then one can construct the BGL-net of standard subspaces on abstract wedge regions. We stress that if the cone $C=\{0\}$, namely there are no non-trivial wedge inclusion, then isotony property is trivially satisfied. Hermitian simple Lie algebras have a non-trivial positive cone and their unitary positive energy representations are studied in [MdR07]. Then if $U$ is a positive energy representation satisfying the spectral condition (HK3) with respect to $C\neq\{0\}$, the isotony property of the BGL-net follows by the inclusion property of standard subspaces, in particular by the converse of the Borchers theorem (cf. Sect. 2.4.2, [MN21, Thm. 4.12]). Once a net of standard subspaces on abstract wedge regions is constructed the second quantization net of von Neumann algebras is canonically provided by relations (12) and (13). If a Lie group $G$ has a non-trivial center, the locality property has to be reformulated in terms of twisted central complements. This has been investigated in [MN21, Sect. 4.2.2]. In this case a proper second quantization procedure respecting a twisted locality conditions has still to be established. Recent results on second quantizations are contained in [CSL23].
@@
Given a unitary representation $(U,\mathcal{H})$ of a connected a Lie group $G$ and a homogeneous space $M=G/H$, we are interested in families $({\sf H}(\mathcal{O}))_{\mathcal{O}\subseteq M}$ of closed real subspaces of $\mathcal{H}$, indexed by open subsets $\mathcal{O}\subseteq M$, called nets of real subspaces on $M$.
@@
    Bisognano–Wichmann property: There exists an open subset $W\subseteq M$ (called a wedge region), such that ${\sf H}(W)$ is standard and $\Delta_{{\sf H}(W)}=e^{2\pi i\partial U(h)}$ for some $h\in{\mathfrak{g}}$.
We remark that locality is not present among the list, since it is to be investigated in [MN25]. Nets satisfying (Iso), (Cov), (RS), (BW) have been constructed on non-compactly causal symmetric spaces in [FNÓ23] and on compactly causal spaces in [NÓ23], using the language of distribution vectors. Given the $G$-orbit of a wedge region, one can define the maximal net on bounded regions by intersection of wedge subspaces ${\sf H}^{\text{max}}(\mathcal{O})=\bigcap\{{\sf H}(gW):\mathcal{O}\subset gW,g\in G\}$, this has been studied in [MN24, Sect. 2.2.4].